OpenStudy (anonymous):

Solve the following system of equations using a algebra method. 3x-5y=-2 y+3y=4

OpenStudy (anonymous):

you could use elimination

OpenStudy (anonymous):

you could do y=mx+b

OpenStudy (anonymous):

you need to get it in that order

OpenStudy (anonymous):

is the second equation \[y+3y =4 ~~or~~x+3y = 4? \]

OpenStudy (anonymous):

this one @Jacaira

OpenStudy (anonymous):

ohhh o.o

OpenStudy (anonymous):

the second one is x+3y=4

OpenStudy (anonymous):

took you forever to reply lol anyway, now you can work on using elimination like i said before

OpenStudy (anonymous):